# Joint semi-supervised and contrastive learning enables zero-shot domain-adaptation and multi-domain segmentation

### v1

@@ IV-F Multi-domain supervised contrastive learning
We train both Baseline and SegCLR models on $D^{s}=D_{\mathrm{All}}\equiv D_{1}\cup D_{2}\cup D_{3}$, evaluating the two trained models on each domain separately. Note that no unlabeled data is used in this setting, where only the images from the annotated training data were used for contrastive learning. Quantitative results are shown in Figure 8 and Table IV, with examples depicted in Figure 9. The results corroborate the findings in [28] in that Baseline trained on $D_{\mathrm{All}}$ outperforms all other individual Baseline models for the corresponding target domains (i.e., , comparing the rows of Baseline across each column), except for a minuscule margin in Dice for $D_{3}$. The observation further extends to SegCLR, i.e., when trained on $D_{\mathrm{All}}$, SegCLR consistently outperforms Baseline as well as all other individual-domain-trained SegCLR results, for all examined domains and both metrics. Evidently, SegCLR framework can effectively leverage contrastive information even from labeled data, thereby enhancing model generalizability across datasets as well as in-domain image variations. Consequently, SegCLR emerges as an ideal framework also for supervised learning only with labeled data, i.e., when there is no domain shift. This further includes settings involving multiple domains for training and/or inference.
TABLE IV: Zero-shot DA results for models trained on different $D^{s}$ (rows) and evaluated on different $D^{t}$ (columns). Absolute metrics are shown averaged across all classes, in bold for the best performance.
$D^{s}$ Model Dice UVD
@@
$D_{\textrm{1}}$ Baseline 67.14$\pm$ 8.24 39.23$\pm$ 8.18 64.77$\pm$ 11.71 6.00$\pm$ 2.32 16.47$\pm$ 2.15 9.38$\pm$ 7.06
SegCLR 69.13$\pm$ 7.94 53.07$\pm$ 8.35 71.13$\pm$ 17.16 5.94$\pm$ 2.45 12.61$\pm$ 2.24 8.27$\pm$ 6.10
@@ V Conclusion
The effectiveness of SegCLR does not rely on the availability of large amounts of unlabeled data, and it performs strikingly well even with limited or no unlabeled data from the target domain, i.e., in unsupervised and zero-shot domain adaptation scenarios. These have been demonstrated in a variety of experimental settings, with the exciting implications as follows: First, on the contrary to what can be anticipated, very large unlabeled datasets are not required to benefit from SegCLR, and even minimal target data is helpful. Second, impressive results can already be attained with no access to or knowledge of the target domain at all! In such zero-shot domain adaptation setting, labeled data is available only from source domain(s), with additional target domains only encountered upon deployment. The versatility of SegCLR is further demonstrated in a multi-domain training experiment where it is concurrently trained on labeled data from multiple domains. The results affirm that SegCLR is not only beneficial for new, unseen domains in a UDA setting, but it also yields superior results in typical supervised settings with labeled data, from both a single or multiple domains. Accordingly, we propose SegCLR to augment conventional supervised UNet, with added generalizability and without compromise.
In summary, our findings indicate that SegCLR provides a powerful and versatile approach to tackle training and inference across multiple domains in medical imaging. With its capability of learning from both labeled and unlabeled data across diverse domains, SegCLR shows great promise for improving the applicability of deep learning models to complex clinical settings involving heterogeneous domains.
@@
UpperBound 67.25$\pm$ 1.66 59.64$\pm$ 3.31 60.39$\pm$ 1.70 57.94$\pm$ 3.02 N/A N/A N/A N/A 11.70$\pm$ 0.36 9.78$\pm$ 0.53 8.19$\pm$ 0.46 8.77$\pm$ 0.46 N/A N/A N/A N/A
Baseline 44.82$\pm$ 4.90 44.22$\pm$ 6.15 34.98$\pm$ 7.72 32.91$\pm$ 6.48 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 18.58$\pm$ 0.92 16.70$\pm$ 1.96 14.45$\pm$ 2.19 16.14$\pm$ 0.98 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 34.78$\pm$ 4.71 35.46$\pm$ 7.18 44.99$\pm$ 9.47 42.32$\pm$ 7.45 52.19$\pm$ 6.97 47.70$\pm$ 3.00 62.03$\pm$ 2.80 47.03$\pm$ 2.97 18.03$\pm$ 1.86 17.68$\pm$ 4.50 15.66$\pm$ 5.97 17.63$\pm$ 9.65 2.98$\pm$ 0.13 11.31$\pm$ 0.54 9.61$\pm$ 0.98 9.25$\pm$ 0.84
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 64.00$\pm$ 6.72 59.16$\pm$ 2.88 56.50$\pm$ 3.98 58.79$\pm$ 5.43 75.51$\pm$ 3.86 62.79$\pm$ 1.70 71.30$\pm$ 1.37 57.97$\pm$ 1.18 12.49$\pm$ 0.80 10.67$\pm$ 1.49 10.99$\pm$ 2.03 12.37$\pm$ 3.20 2.63$\pm$ 0.13 8.64$\pm$ 0.49 5.56$\pm$ 0.28 7.06$\pm$ 0.42
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 61.29$\pm$ 7.87 58.30$\pm$ 3.06 57.09$\pm$ 4.06 56.62$\pm$ 4.57 76.22$\pm$ 5.23 63.20$\pm$ 1.84 71.31$\pm$ 2.11 61.01$\pm$ 2.12 13.07$\pm$ 1.75 10.76$\pm$ 2.12 10.96$\pm$ 2.45 11.70$\pm$ 2.79 2.52$\pm$ 0.12 8.50$\pm$ 0.58 5.40$\pm$ 0.23 6.96$\pm$ 0.41
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 52.76$\pm$ 10.43 46.87$\pm$ 10.14 51.74$\pm$ 3.14 51.03$\pm$ 10.30 74.25$\pm$ 2.87 62.46$\pm$ 0.89 71.44$\pm$ 0.82 57.36$\pm$ 1.32 14.76$\pm$ 4.17 15.02$\pm$ 4.66 11.48$\pm$ 2.55 11.74$\pm$ 1.35 2.33$\pm$ 0.08 8.50$\pm$ 0.29 6.02$\pm$ 0.31 7.66$\pm$ 0.26
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 62.44$\pm$ 4.73 59.30$\pm$ 6.18 55.86$\pm$ 2.05 54.78$\pm$ 6.01 75.85$\pm$ 3.82 63.79$\pm$ 1.51 69.67$\pm$ 1.96 59.86$\pm$ 2.80 13.20$\pm$ 1.81 10.75$\pm$ 1.49 11.12$\pm$ 2.30 11.17$\pm$ 1.69 2.60$\pm$ 0.17 8.81$\pm$ 0.64 5.52$\pm$ 0.39 7.37$\pm$ 0.49
@@
UpperBound 65.45$\pm$ 1.09 93.18$\pm$ 0.85 N/A N/A N/A N/A 11.36$\pm$ 0.23 1.79$\pm$ 0.17 N/A N/A N/A N/A
Baseline 53.84$\pm$ 3.26 75.71$\pm$ 3.64 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 16.23$\pm$ 0.99 2.53$\pm$ 0.24 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 50.46$\pm$ 2.64 75.79$\pm$ 12.31 67.74$\pm$ 6.05 61.00$\pm$ 2.09 70.09$\pm$ 1.49 53.19$\pm$ 1.62 17.24$\pm$ 0.56 3.75$\pm$ 2.09 2.55$\pm$ 0.11 8.54$\pm$ 0.84 6.05$\pm$ 0.24 7.85$\pm$ 0.56
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 56.45$\pm$ 4.12 86.51$\pm$ 4.18 78.03$\pm$ 2.16 64.48$\pm$ 2.08 71.06$\pm$ 1.62 60.98$\pm$ 2.91 14.12$\pm$ 0.85 2.79$\pm$ 0.71 2.50$\pm$ 0.09 8.55$\pm$ 0.84 5.33$\pm$ 0.46 6.94$\pm$ 0.55
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 58.15$\pm$ 3.17 88.89$\pm$ 3.21 77.82$\pm$ 2.38 63.14$\pm$ 1.86 70.92$\pm$ 2.14 61.11$\pm$ 2.04 13.59$\pm$ 0.76 2.10$\pm$ 0.54 2.51$\pm$ 0.12 8.83$\pm$ 0.61 5.23$\pm$ 0.18 7.16$\pm$ 0.47
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 56.17$\pm$ 2.47 73.84$\pm$ 8.96 73.78$\pm$ 3.06 61.66$\pm$ 2.36 70.74$\pm$ 0.95 57.21$\pm$ 2.02 16.30$\pm$ 0.69 3.33$\pm$ 0.63 2.32$\pm$ 0.08 8.81$\pm$ 0.29 6.06$\pm$ 0.54 7.62$\pm$ 0.37
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 58.39$\pm$ 3.46 89.60$\pm$ 3.86 79.10$\pm$ 2.48 63.72$\pm$ 1.71 70.98$\pm$ 1.84 60.88$\pm$ 2.53 13.13$\pm$ 0.62 1.99$\pm$ 0.68 2.58$\pm$ 0.20 8.32$\pm$ 0.60 5.36$\pm$ 0.59 7.10$\pm$ 0.66
TABLE S4: Zero-shot DA results for models evaluated on $D_{1}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice UVD
@@
$D_{\textrm{1}}$ Baseline 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 2.26$\pm$ 0.11 8.15$\pm$ 0.30 6.22$\pm$ 0.47 7.38$\pm$ 0.27
SegCLR 80.51$\pm$ 1.89 64.25$\pm$ 1.55 71.35$\pm$ 1.15 60.43$\pm$ 2.43 2.43$\pm$ 0.10 8.89$\pm$ 0.66 5.31$\pm$ 0.21 7.14$\pm$ 0.50
@@
SegCLR 79.84$\pm$ 3.07 64.05$\pm$ 2.18 72.42$\pm$ 1.53 63.14$\pm$ 2.31 2.35$\pm$ 0.09 8.54$\pm$ 0.53 5.01$\pm$ 0.21 7.14$\pm$ 0.53
TABLE S5: Zero-shot DA results for models evaluated on $D_{2}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice UVD
@@
$D_{\textrm{1}}$ Baseline 44.82$\pm$ 4.90 44.22$\pm$ 6.15 34.98$\pm$ 7.72 32.91$\pm$ 6.48 18.58$\pm$ 0.92 16.70$\pm$ 1.96 14.45$\pm$ 2.19 16.14$\pm$ 0.98
SegCLR 59.66$\pm$ 7.82 52.37$\pm$ 10.61 49.93$\pm$ 3.99 50.33$\pm$ 6.67 13.77$\pm$ 1.54 12.21$\pm$ 2.92 12.54$\pm$ 2.27 11.94$\pm$ 1.86
@@
SegCLR 74.27$\pm$ 1.92 67.80$\pm$ 4.46 71.05$\pm$ 2.72 64.80$\pm$ 3.22 9.58$\pm$ 0.70 8.66$\pm$ 0.38 6.98$\pm$ 0.69 7.52$\pm$ 0.37
TABLE S6: Zero-shot DA results for models evaluated on $D_{3}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes. Note that PED and SHRM are not relevant for $D_{3}$ (see Section III).
$D^{s}$ Model Dice UVD
@@
$D_{\textrm{1}}$ Baseline 53.84$\pm$ 3.26 75.71$\pm$ 3.64 16.23$\pm$ 0.99 2.53$\pm$ 0.24
SegCLR 54.69$\pm$ 2.88 87.58$\pm$ 3.58 14.17$\pm$ 0.79 2.37$\pm$ 0.66

### v3

@@ IV-F Multi-domain supervised contrastive learning
We train both UNet and SegCLR models on $D^{s}=D_{\mathrm{All}}\equiv D_{1}\cup D_{2}\cup D_{3}$, evaluating the two trained models on each domain separately. Note that no unlabeled data is used in this setting, where only the images from the annotated training data were used for contrastive learning. Quantitative results are shown in Figure 8 and Table IV, with examples depicted in Figure 9. The results corroborate the findings in [33] in that UNet trained on $D_{\mathrm{All}}$ outperforms all other individual UNet models for the corresponding target domains (i.e., comparing the rows of UNet across each column), except for a minuscule margin in Dice for $D_{3}$. The observation further extends to SegCLR, i.e., when trained on $D_{\mathrm{All}}$, SegCLR consistently outperforms UNet as well as all other individual-domain-trained SegCLR results, for all examined domains and both metrics. Evidently, SegCLR framework can effectively leverage contrastive information even from labeled data, thereby enhancing model generalizability across datasets as well as in-domain image variations. Consequently, SegCLR emerges as an ideal framework also for supervised learning only with labeled data, i.e., when there is no domain shift. This further includes settings involving multiple domains for training and/or inference.
TABLE IV: Domain generalization results for models trained on different $D^{s}$ (rows) and evaluated on different $D^{t}$ (columns). Absolute metrics are shown averaged across all classes, in bold for the best performance.
$D^{s}$ Model Dice [%] UVD [fL]
@@
$D_{\textrm{1}}$ UNet 67.14$\pm$ 8.24 39.23$\pm$ 8.18 64.77$\pm$ 11.71 600$\pm$ 232 1647$\pm$ 215 938$\pm$ 706
SegCLR 69.13$\pm$ 7.94 53.07$\pm$ 8.35 71.13$\pm$ 17.16 594$\pm$ 245 1261$\pm$ 224 827$\pm$ 610
@@ V Conclusion
The effectiveness of SegCLR does not rely on the availability of large amounts of unlabeled data, and it performs strikingly well even with limited or no unlabeled data from the target domain, i.e., in unsupervised domain adaptation and domain generalization scenarios. These have been demonstrated in a variety of experimental settings, with the exciting implications as follows: First, on the contrary to what can be anticipated, very large unlabeled datasets are not required to benefit from SegCLR, and even minimal target data is helpful. Second, impressive results can already be attained with no access to or knowledge of the target domain at all! In such domain generalization setting, labeled data is available only from source domain(s), with additional target domains only encountered upon deployment.
The versatility of SegCLR is further demonstrated in a multi-domain training experiment where it is concurrently trained on labeled data from multiple domains. The results affirm that SegCLR is not only beneficial for new, unseen domains in a UDA setting, but it also yields superior results in typical supervised settings with labeled data, from both a single or multiple domains. Accordingly, we propose SegCLR to augment conventional supervised UNet, with added generalizability and without compromise.
In summary, our findings indicate that SegCLR provides a powerful and versatile approach to tackle training and inference across multiple domains in medical imaging. With its capability of learning from both labeled and unlabeled data across diverse domains, SegCLR shows great promise for improving the applicability of deep learning models to complex clinical settings involving heterogeneous domains. Furthermore, in contrast to other UDA approaches, SegCLR can retain source domain knowledge while adapting to target domains, which can facilitate its use and adaptation for future applications in continual and incremental learning.
@@
TABLE S2: Datasets employed as source ($D^{s}$), target ($D^{t}$) in the experimental setting for each of the paper sections.
Section & experiment Training Evaluation
Labeled ($D^{s}$) Unlabeled ($D^{t}$)
4.1. Cross-device DA $D_{1}$ $D_{2}$ $D_{1}$; $D_{2}$
4.2. Cross-disease DA $D_{1}$ $D_{3}$ $D_{1}$; $D_{3}$
4.4. Reducing data $D_{1}$ $D_{2}$ $D_{1}$; $D_{2}$
4.5. Domain generalization $D_{1}$; $D_{2}$; $D_{3}$; $D_{1}\!\cup\!D_{2}\!\cup\!D_{3}$ N/A $D_{1}$; $D_{2}$; $D_{3}$; $D_{1}\!\cup\!D_{2}\!\cup\!D_{3}$
@@ 4.4. Reducing data $D_{1}$ $D_{2}$ $D_{1}$; $D_{2}$
UNet (UpperBound) 67.25$\pm$ 1.66 59.64$\pm$ 3.31 60.39$\pm$ 1.70 57.94$\pm$ 3.02 N/A N/A N/A N/A 1170$\pm$ 36 978$\pm$ 53 819$\pm$ 46 877$\pm$ 46 N/A N/A N/A N/A
UNet (Baseline) 44.82$\pm$ 4.90 44.22$\pm$ 6.15 34.98$\pm$ 7.72 32.91$\pm$ 6.48 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 1858$\pm$ 92 1670$\pm$ 196 1445$\pm$ 219 1614$\pm$ 98 226$\pm$ 11 815$\pm$ 30 622$\pm$ 47 738$\pm$ 27
@@
SimSiam 45.37$\pm$ 3.83 48.94$\pm$ 4.52 36.09$\pm$ 5.45 34.17$\pm$ 5.72 78.81$\pm$ 1.94 63.26$\pm$ 1.23 71.15$\pm$ 0.77 56.37$\pm$ 1.46 1800$\pm$ 98 1630$\pm$ 149 1395$\pm$ 127 1566$\pm$ 118 222$\pm$ 7 854$\pm$ 43 595$\pm$ 44 742$\pm$ 37
[…]
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 34.78$\pm$ 4.71 35.46$\pm$ 7.18 44.99$\pm$ 9.47 42.32$\pm$ 7.45 52.19$\pm$ 6.97 47.70$\pm$ 3.00 62.03$\pm$ 2.80 47.03$\pm$ 2.97 1803$\pm$ 186 1768$\pm$ 450 1566$\pm$ 597 1763$\pm$ 965 298$\pm$ 13 1131$\pm$ 54 961$\pm$ 98 925$\pm$ 84
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 64.00$\pm$ 6.72 59.16$\pm$ 2.88 56.50$\pm$ 3.98 58.79$\pm$ 5.43 75.51$\pm$ 3.86 62.79$\pm$ 1.70 71.30$\pm$ 1.37 57.97$\pm$ 1.18 1249$\pm$ 80 1067$\pm$ 149 1099$\pm$ 203 1237$\pm$ 320 263$\pm$ 13 864$\pm$ 49 556$\pm$ 28 706$\pm$ 42
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 61.29$\pm$ 7.87 58.30$\pm$ 3.06 57.09$\pm$ 4.06 56.62$\pm$ 4.57 76.22$\pm$ 5.23 63.20$\pm$ 1.84 71.31$\pm$ 2.11 61.01$\pm$ 2.12 1307$\pm$ 175 1076$\pm$ 212 1096$\pm$ 245 1170$\pm$ 279 252$\pm$ 12 850$\pm$ 58 540$\pm$ 23 696$\pm$ 41
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 52.76$\pm$ 10.43 46.87$\pm$ 10.14 51.74$\pm$ 3.14 51.03$\pm$ 10.30 74.25$\pm$ 2.87 62.46$\pm$ 0.89 71.44$\pm$ 0.82 57.36$\pm$ 1.32 1476$\pm$ 417 1502$\pm$ 466 1148$\pm$ 255 1174$\pm$ 135 233$\pm$ 8 850$\pm$ 29 602$\pm$ 31 766$\pm$ 26
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 62.44$\pm$ 4.73 59.30$\pm$ 6.18 55.86$\pm$ 2.05 54.78$\pm$ 6.01 75.85$\pm$ 3.82 63.79$\pm$ 1.51 69.67$\pm$ 1.96 59.86$\pm$ 2.80 1320$\pm$ 181 1075$\pm$ 149 1112$\pm$ 230 1117$\pm$ 169 260$\pm$ 17 881$\pm$ 64 552$\pm$ 39 737$\pm$ 49
@@
UNet (UpperBound) 65.45$\pm$ 1.09 93.18$\pm$ 0.85 N/A N/A N/A N/A 1136$\pm$ 23 179$\pm$ 17 N/A N/A N/A N/A
UNet (Baseline) 53.84$\pm$ 3.26 75.71$\pm$ 3.64 77.72$\pm$ 2.07 63.64$\pm$ 1.13 70.98$\pm$ 0.57 56.20$\pm$ 0.99 1623$\pm$ 99 253$\pm$ 24 226$\pm$ 11 815$\pm$ 30 622$\pm$ 47 738$\pm$ 27
@@
SimSiam 54.71$\pm$ 3.69 76.70$\pm$ 5.77 76.46$\pm$ 2.89 63.21$\pm$ 1.18 71.18$\pm$ 0.88 56.76$\pm$ 1.61 1588$\pm$ 82 254$\pm$ 41 230$\pm$ 12 831$\pm$ 32 583$\pm$ 37 737$\pm$ 26
@@
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{pool}}$) 50.46$\pm$ 2.64 75.79$\pm$ 12.31 67.74$\pm$ 6.05 61.00$\pm$ 2.09 70.09$\pm$ 1.49 53.19$\pm$ 1.62 1724$\pm$ 56 375$\pm$ 209 255$\pm$ 11 854$\pm$ 84 605$\pm$ 24 785$\pm$ 56
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{pool}}$) 56.45$\pm$ 4.12 86.51$\pm$ 4.18 78.03$\pm$ 2.16 64.48$\pm$ 2.08 71.06$\pm$ 1.62 60.98$\pm$ 2.91 1412$\pm$ 85 279$\pm$ 71 250$\pm$ 9 855$\pm$ 84 533$\pm$ 46 694$\pm$ 55
SegCLR($P_{\mathrm{a}}$,$C_{\mathrm{ch}}$) 58.15$\pm$ 3.17 88.89$\pm$ 3.21 77.82$\pm$ 2.38 63.14$\pm$ 1.86 70.92$\pm$ 2.14 61.11$\pm$ 2.04 1359$\pm$ 76 210$\pm$ 54 251$\pm$ 12 883$\pm$ 61 523$\pm$ 18 716$\pm$ 47
SegCLR($P_{\mathrm{s}}$,$C_{\mathrm{ch}}$) 56.17$\pm$ 2.47 73.84$\pm$ 8.96 73.78$\pm$ 3.06 61.66$\pm$ 2.36 70.74$\pm$ 0.95 57.21$\pm$ 2.02 1630$\pm$ 69 333$\pm$ 63 232$\pm$ 8 881$\pm$ 29 606$\pm$ 54 762$\pm$ 37
SegCLR($P_{\mathrm{s+a}}$,$C_{\mathrm{ch}}$) 58.39$\pm$ 3.46 89.60$\pm$ 3.86 79.10$\pm$ 2.48 63.72$\pm$ 1.71 70.98$\pm$ 1.84 60.88$\pm$ 2.53 1313$\pm$ 62 199$\pm$ 68 258$\pm$ 20 832$\pm$ 60 536$\pm$ 59 710$\pm$ 66
TABLE S5: Domain generalization results for models evaluated on $D_{1}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice [%] UVD [fL]
@@
SegCLR 41.58$\pm$ 6.57 38.63$\pm$ 3.21 N/ A N/ A 857$\pm$ 356 1342$\pm$ 79 N/ A N/ A
@@
SegCLR 79.84$\pm$ 3.07 64.05$\pm$ 2.18 72.42$\pm$ 1.53 63.14$\pm$ 2.31 235$\pm$ 9 854$\pm$ 53 501$\pm$ 21 714$\pm$ 53
TABLE S6: Domain generalization results for models evaluated on $D_{2}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes.
$D^{s}$ Model Dice [%] UVD [fL]
@@
SegCLR 45.37$\pm$ 6.81 35.62$\pm$ 9.37 N/ A N/ A 2898$\pm$ 1293 2127$\pm$ 530 N/ A N/ A
@@
SegCLR 74.27$\pm$ 1.92 67.80$\pm$ 4.46 71.05$\pm$ 2.72 64.80$\pm$ 3.22 958$\pm$ 70 866$\pm$ 38 698$\pm$ 69 752$\pm$ 37
TABLE S7: Domain generalization results for models evaluated on $D_{3}$ after training on different $D^{s}$. Absolute metrics are shown for each of the evaluated classes. Note that PED and SHRM are not relevant for $D_{3}$ (see Section III).
$D^{s}$ Model Dice [%] UVD [fL]